OpenStudy (anonymous):

Find the thirty-second term of the following sequence. 9, 15, 21, ...

OpenStudy (anonymous):

186 199 195

OpenStudy (saifoo.khan):

Which one did you got?

OpenStudy (anonymous):

i got 201

OpenStudy (anonymous):

whats the formula

OpenStudy (saifoo.khan):

It's an arithmetic series. difference is 6. Formula, t_n = a + (n-1)d t_34 = 9 + (34-1)6

OpenStudy (anonymous):

o still dont get an answer from above

OpenStudy (anonymous):

Saifoo.khan provided us with an equation that "models" your question. That is, you would sub in the "x"th term that you want to find out in the sequence, and you get it. Suppose you want t_x = t_34 Then, you will get: t_34 = 9 + (34-1)*6

OpenStudy (saifoo.khan):

hey the term is 32nd. so 9 + (32-1)6

OpenStudy (saifoo.khan):

i did a mistake.

OpenStudy (anonymous):

195

OpenStudy (saifoo.khan):

\[\checkmark \]

OpenStudy (anonymous):

ok thank you @saifoo.khan

OpenStudy (saifoo.khan):

No problem.